OpenStudy (anonymous):

Form a polynomial whose zeros and degrees are given : zeros:-2,2,9 degree:3

OpenStudy (anonymous):

Expand (x+2)(x-2)(x-9) = (x²-4)(x-9) = x^3-9x²-4x+36